# Generative Autoregressive Transformers for Model-Agnostic Federated MRI Reconstruction

### v1

@@ 6 Discussion
Devised to accurately capture an underlying distribution, generative priors synthesize MR images depicting random anatomical structures that do not correspond to the precise anatomy of any actual subject [75, 80]. In FedGAT, such synthetic MR images from separate sites are employed in training of site-specific reconstruction models to improve generalization across sites. Our assessments based on FID scores indicate that the synthetic MR images generated by the GAT prior attain a level of quality and realism closely mimicking actual MR images. Furthermore, reconstruction models trained on these synthetic MR images yield high performance when independently evaluated on actual MRI acquisitions in the test sets. Still, in the absence of anatomical guidance, the stochastic nature of generative priors naturally induce a degree of susceptibility to artificial features in synthetic MR images [81]. Our observations indicate that the downstream reconstruction models trained on synthetic MR images in this study are not unduly influenced by undesirable hallucinations. When necessary, hallucinations might be alleviated by building reconstruction models equipped with physics-driven modules and learning objectives [82], or building generative priors via architectures that offer enhanced capture of contextual relationships among tissue signals [83, 84]. It remains important future work to validate the proposed method and its anatomical fidelity on broader patient cohorts.

### v2

@@ 6.1 Site Specialization
6.2 Computational Considerations
FedGAT might offer a unique set of computational advantages for building MRI reconstruction models compared to existing FL frameworks. In conventional FL, sites must agree on a reconstruction model of common architecture, and each communication round involves training and aggregating local copies of this model. If a different architecture is later preferred, the entire FL training must be repeated, incurring substantial computational costs for each imaging site [30]. FedGAT avoids this overhead by training a shared generative prior to model the distribution of multi-site MRI data independently of reconstruction. Once trained, the prior can be reused to synthesize data for locally building reconstruction models of various architectures, amortizing the cost of FL training across experiments on multiple architectures. In distillation-based FL methods, model heterogeneity is supported by maintaining separate site-specific and global reconstruction models, and aligning between these models through repeated distillation steps on the server and/or at individual sites [75]. FedGAT bypasses this overhead by enabling cross-site knowledge transfer through synthetic image generation, eliminating the need for direct model alignment. While previous generative FL methods also adopt a decoupled training approach based on adversarial or diffusion priors [59], FedGAT offers a refined balance across sample fidelity and inference cost. Our results indicate that the GAT prior achieves higher image fidelity than both adversarial and diffusion priors, while requiring moderate inference time—faster than diffusion (10 autoregressive stages vs. 1000 diffusion steps) and slower than adversarial methods (which require just one step).
6.3 Fidelity of Synthetic Images
Devised to capture an underlying distribution, generative priors synthesize MR images depicting random anatomical structures that do not correspond to the precise anatomy of any actual subject [87, 92]. In FedGAT, such synthetic MR images from separate sites are employed in training of site-specific reconstruction models to improve generalization across sites. Our assessments based on FID scores indicate that the synthetic MR images generated by the GAT prior attain a level of quality and diversity closely mimicking actual MR images. Furthermore, reconstruction models trained on these synthetic MR images yield high performance when independently evaluated on actual MRI acquisitions in the test sets. Still, in the absence of anatomical guidance, the stochastic nature of generative priors naturally induce a degree of susceptibility to artificial features in synthetic MR images [93]. Our observations indicate that the downstream reconstruction models trained on synthetic MR images in this study are not unduly influenced by undesirable hallucinations. When necessary, hallucinations might be alleviated by building reconstruction models equipped with physics-driven modules and learning objectives [94], or building generative priors via architectures that offer enhanced capture of contextual relationships among tissue signals [95, 96]. It remains important future work to validate the proposed method and its anatomical fidelity on broader patient cohorts.
6.4 Patient Privacy